# Sparse Bayesian multidimensional scaling(s)

### v1

@@ Abstract
Bayesian multidimensional scaling (BMDS) is a probabilistic dimension reduction tool that allows one to model and visualize data consisting of dissimilarities between pairs of objects. Although BMDS has proven useful within, e.g., Bayesian phylogenetic inference, its likelihood and gradient calculations require a burdensome $\mathcal{O}(N^{2})$ floating-point operations, where $N$ is the number of data points. Thus, BMDS becomes impractical as $N$ grows large. We propose and compare two sparse versions of BMDS (sBMDS) that apply log-likelihood and gradient computations to subsets of the observed dissimilarity matrix data. Landmark sBMDS (L-sBMDS) extracts columns, while banded sBMDS (B-sBMDS) extracts diagonals of the data. These sparse variants let one specify a time complexity between $N^{2}$ and $N$. Under simplified settings, we prove posterior consistency for subsampled distance matrices. Through simulations, we examine the accuracy and computational efficiency across all models using both the Metropolis-Hastings and Hamiltonian Monte Carlo algorithms. We observe approximately 3-fold, 10-fold and 40-fold speedups with negligible loss of accuracy, when applying the sBMDS likelihoods and gradients to 500, 1,000 and 5,000 data points with 50 bands (landmarks); these speedups only increase with the size of data considered. Finally, we apply the sBMDS variants to the phylogeographic modeling of multiple influenza subtypes to better understand how these strains spread through global air transportation networks.
@@ 1 Introduction
Multidimensional scaling (MDS) is a dimension reduction technique that maps pairwise dissimilarity measurements corresponding to a set of $N$ objects to a configuration of $N$ points within a low-dimensional Euclidean space (Torgerson,, 1952). Classical MDS uses the spectral decomposition of a doubly centered matrix derived from the observed dissimilarity matrix to calculate the objects’ coordinates. While classical MDS serves as a valuable data visualization tool, probabilistic extensions further enable uncertainty quantification in the context of Bayesian hierarchical models. Oh and Raftery, (2001) propose a Bayesian framework for MDS (BMDS) under the assumption that the observed dissimilarities follow independent truncated normal probability density functions (PDFs). BMDS facilitates Bayesian inference of object configurations in a manner that is robust to violations of the Euclidean model assumption and dimension misspecifications. Thanks to its probabilistic nature, one may integrate BMDS into hierarchical modeling approaches for Bayesian phylogeography (Bedford et al.,, 2014; Holbrook et al.,, 2021; Li et al.,, 2023), clustering (Man-Suk and Raftery,, 2007), and variable selection (Lin and Fong,, 2019).
Unfortunately, BMDS is difficult to scale to big data settings; computing the BMDS log-likelihood and gradient each have $\mathcal{O}(N^{2})$ complexity. Bedford et al., (2014) partially circumvent this problem by assuming that the observed data follow non-truncated Gaussian distributions, thereby avoiding the costly floating-point operations necessary to evaluate the Gaussian cumulative density function (CDF) in the truncated normal PDFs (2). However, this method does not hold for non-negative quantities, leading to inference from an incorrect model. Holbrook et al., (2021) mitigate BMDS’s computational burden through massive parallelization using multi-core central processing units, vectorization and graphic processing units. They obtain substantial performance gains, but parallelization requires expensive hardware. In either case, these models still scale quadratically in the number of objects. We therefore develop a framework that reduces the time complexity to $\mathcal{O}(N)$ by inducing sparsity on the observed dissimilarity matrix.
@@
We perform experiments with simulated data and show that our sparse versions of BMDS (sBMDS) obtain significant speedups while preserving inferential accuracy. We also use sBMDS to analyze the geographic spread of four influenza subtypes. Bayesian phylogeography studies how viruses, bacteria, or pathogens evolve and interact over time and location. Bedford et al., (2014) simultaneously characterize antigenic and genetic patterns of influenza by combining BMDS with an evolutionary diffusion process on the latent strain locations. They apply BMDS on hemmagglutination inhibition assay data to place the subtypes on a low-dimensional antigenic map. Holbrook et al., (2021) implement a similar Bayesian phylogenetic MDS model, but perform phylogeographic inference on pairwise distances arising from air traffic data. Additionally, Li et al., (2023) use phylogenetic BMDS on pairwise distances stemming from hepaciviruses to infer the viral locations in a lower dimensional geographic and host space. We extend these types of models under sparse assumptions.
@@
In the following, we present two versions of sparse BMDS and prove that under simplistic conditions, the posterior latent locations are consistent for subsampled distance matrices (Section 2). In Section 3, we evaluate the empirical accuracy and computational performance of both methods. We then apply sBMDS to the phylogeographic modeling of influenza variants and verify that we obtain similar migration rate estimates for both full and sparse BMDS models (Section 3.2). We conclude by summarizing our findings and discussing future research directions (Section 4).
@@ 2.1 Bayesian multidimensional scaling
Bayesian multidimensional scaling (BMDS) (Oh and Raftery,, 2001) models a set of $N$ objects’ locations as latent variables in low-dimensional space under the assumption that the observed dissimilarity measures follow a prescribed joint probability distribution. Let $N_{A}(\mu,\sigma^{2})$ denote the Gaussian distribution truncated to $A$ for $A\subset\mathbb{R}$. Within BMDS, each observed dissimilarity measure $\delta_{nn^{\prime}}$ is the posited latent measure $\delta_{nn^{\prime}}^{*}$ plus a truncated Gaussian error:
$\delta_{nn^{\prime}}\sim N_{(0,\infty)}(\delta_{nn^{\prime}}^{*},\sigma^{2}),\ n\neq n^{\prime},\ n,n^{\prime}=[N]=\{1,\dots,N\},$ (1)
@@
$\ell(\bm{\Delta},\sigma^{2})\propto-\frac{m}{2}\log(\sigma^{2})-\sum_{n>n^{\prime}}\biggr{[}\frac{(\delta_{nn^{\prime}}-\delta_{nn^{\prime}}^{*})^{2}}{2\sigma^{2}}+\log\Phi\biggr{(}\frac{\delta_{nn^{\prime}}^{*}}{\sigma}\biggr{)}\biggr{]},$ (2)
@@
$\nabla{\mathbf{x}_{n}}\ell(\bm{\Delta},\sigma^{2})=-\sum_{n\neq n^{\prime}}\biggr{[}\frac{(\delta_{nn^{\prime}}^{*}-\delta_{nn^{\prime}})}{\sigma^{2}}+\frac{\phi(\delta_{nn^{\prime}}^{*}/\sigma)}{\sigma\Phi(\delta_{nn^{\prime}}^{*}/\sigma)}\ \frac{(\mathbf{x}_{n}-\mathbf{x}_{n^{\prime}})}{\delta_{nn^{\prime}}^{*}}\biggr{]}\equiv-\sum_{n\neq n^{\prime}}\mathbf{r}_{nn^{\prime}}.$ (3)
@@ 2.2 Sparse likelihoods and their gradients
$\ell(\bm{\Delta},\sigma^{2})\propto-\frac{m^{*}}{2}\log(\sigma^{2})-\sum_{n=1}^{N}\sum_{n^{\prime}\in J_{n,N}}\biggr{[}\frac{(\delta_{nn^{\prime}}-\delta_{nn^{\prime}}^{*})^{2}}{2\sigma^{2}}+\log\Phi\biggr{(}\frac{\delta_{nn^{\prime}}^{*}}{\sigma}\biggr{)}\biggr{]},$ (4)
@@
$\nabla{\mathbf{x}_{n}}\ell(\bm{\Delta},\sigma^{2})=-\sum_{n=1}^{N}\sum_{n^{\prime}\in J_{n,N}}\mathbf{r}_{nn^{\prime}}.$ (5)
We reduce the computational complexity of BMDS by including a small subset of couplings $J_{n,N}$ per object $n$, where $|J_{n,N}|\ll N$, and $m^{*}=\sum_{n=1}^{N}|J_{n,N}|$ is the number of couplings. Here, we discuss two possible strategies for choosing $J_{n,N}$. The first option is to extract $B\in[N-1]$ off-diagonal bands of the observed distance matrix such that $J_{n,N}=[\max(1,n-B),\min(N,n+B)]\backslash\{n\}$ for all $n$. The second approach is to choose $L\in[N]$ objects called “landmarks” and select each landmark’s dissimilarities from the remaining $N-1$ objects, e.g., $J_{n,N}=[N]\backslash\{n\}$ for $n\in[L]$ and $J_{n,N}=[L]$ for $n\not\in[L]$. Essentially, this strategy retains a rectangular subset of the observed distance matrix by extracting $L$ columns (rows) of the data. We refer to the first method as banded sBMDS (B-sBMDS) and the second as landmark sBMDS (L-sBMDS).
@@
Table 1: We extract the $(n,n^{\prime})$ pair from the off-diagonals of the observed and latent dissimilarity matrices for banded sBMDS (B-sBMDS) versus the columns for landmark sBMDS (L-sBMDS) to compute their respective log-likelihoods. The $+$ symbol indicates all couplings above are also included. The table on the right shows the calculated log-likelihoods as a function of the number of bands/landmarks. Importantly, the bottom log-likelihoods are equal for both sBMDS variants and correspond to the full BMDS log-likelihood.
@@
Table 2: We extract the $(n,n^{\prime})$ pair from the observed and latent dissimilarity matrices to calculate the sBMDS gradients. On the left, the -- symbol, as in $(n,a-c)$, indicates pairs $(n,a);(n,b);(n,c)$. For example, pair $(3;4-5)$ means we include both pair $(3,4)$ and $(3,5)$. The $+$ symbol indicates all couplings to the left are also included. On the right is the gradient computed for each $\mathbf{x}_{n}$ of $\mathbf{X}$ as function of the number of bands/landmarks. Extracting the entire column of a landmark point gives the full BMDS gradient in $\mathbb{R}^{D}$ whereas banded sBMDS incrementally adds information to the row-wise gradients. Importantly, the rightmost gradients are equal for both sBMDS variants and correspond to the full BMDS gradient.
For B-sBMDS, the number of couplings is the number of elements in $B$ bands. The relationship between the number of bands and number of couplings $C$ is $C=\sum_{b=1}^{B}(N-b)$. We add one less coupling for each additional band. When the number of bands equals $N-1$, we return to the full BMDS case. Using a subset of the observed distance matrix reduces the burden of computing the BMDS likelihood and gradient to $\mathcal{O}(NB)$. Similar arguments hold for L-sBMDS, the likelihoods and gradients of which exhibit $\mathcal{O}(NL)$ time complexity.
For classical MDS, an analogous strategy to L-sBMDS already exists. In MDS, the rate limiting step is the calculation of the top $D$ eigenvalues and eigenvectors from a $N\times N$ matrix. Silva and Tenenbaum, (2004) propose applying classical MDS to $L$ landmark points, e.g., an $L\times N$ submatrix of the observed distance matrix, and then following a distance-based triangulation procedure to determine the remaining object coordinates. L-sBMDS uses the concept of randomly selecting $L$ landmarks as well, but integrates them into the BMDS framework, allowing inference on the entire model. Raftery et al., (2012) approximate the likelihood of their network data by taking a random subset of objects deemed to have no link, reducing the time complexity from $\mathcal{O}(N^{2})$ to $\mathcal{O}(N)$. In the context of a very different network model, they incorporate an array of covariates to model the probability of a link between objects $n$ and $n^{\prime}$ while our model is simpler, using no outside information to aid in determining locations in a latent space.

### v3

@@ Abstract
Bayesian multidimensional scaling (BMDS) is a probabilistic dimension reduction tool that allows one to model and visualize data consisting of dissimilarities between pairs of objects. Although BMDS has proven useful within, e.g., Bayesian phylogenetic inference, its likelihood and gradient calculations require a burdensome $\mathcal{O}(N^{2})$ floating-point operations, where $N$ is the number of data points. Thus, BMDS becomes impractical as $N$ grows large. We propose and compare two sparse versions of BMDS (sBMDS) that apply log-likelihood and gradient computations to subsets of the observed dissimilarity matrix data. Landmark sBMDS (L-sBMDS) extracts columns, while banded sBMDS (B-sBMDS) extracts diagonals of the data. These sparse variants let one specify a time complexity between $N^{2}$ and $N$. Under simplified settings, we prove posterior consistency for subsampled distance matrices. Through simulations, we examine the accuracy and computational efficiency across all models using both the Metropolis-Hastings and Hamiltonian Monte Carlo algorithms. We observe approximately 3-fold, 10-fold and 40-fold speedups with negligible loss of accuracy, when applying the sBMDS likelihoods and gradients to 500, 1,000 and 5,000 data points with 50 bands (landmarks); these speedups only increase with the size of data considered. Finally, we apply the sBMDS variants to: 1) the phylogeographic modeling of multiple influenza subtypes to better understand how these strains spread through global air transportation networks and 2) the clustering of ArXiv manuscripts based on low-dimensional representations of article abstracts. In the first application, sBMDS contributes to holistic uncertainty quantification within a larger Bayesian hierarchical model. In the second, sBMDS provides uncertainty quantification for a downstream modeling task.
@@ 1 Introduction
Multidimensional scaling (MDS) is a dimension reduction technique that maps pairwise dissimilarity measurements corresponding to a set of $N$ objects to a configuration of $N$ points within a low-dimensional Euclidean space (Torgerson,, 1952). Classical MDS uses the spectral decomposition of a doubly centered matrix derived from the observed dissimilarity matrix to calculate the objects’ coordinates. While classical MDS serves as a valuable data visualization tool, probabilistic extensions further enable uncertainty quantification in the context of Bayesian hierarchical models. Oh and Raftery, (2001) propose a Bayesian framework for MDS (BMDS) under the assumption that the observed dissimilarities follow independent truncated normal probability density functions (PDFs). BMDS facilitates Bayesian inference of object configurations in a manner that is robust to violations of the Euclidean model assumption and dimension misspecifications (Oh and Raftery,, 2001). The key benefits of a Bayesian approach to MDS are that it provides uncertainty quantification for the projection itself and conditional distributions that can be easily integrated with other probability models, enabling fully model-based approaches to analyzing dissimilarity data. For example, one may incorporate BMDS into hierarchical modeling frameworks for Bayesian phylogeography (Bedford et al.,, 2014; Holbrook et al.,, 2021; Li et al.,, 2023), clustering (Man-Suk and Raftery,, 2007), and variable selection (Lin and Fong,, 2019).
@@
Bayesian phylogeography uses molecular data from species such as viruses, bacteria or pathogens to probabilistically model their evolution over both time and space (Lemey et al.,, 2009). For instance, one can reconstruct viral dispersion patterns to better understand the way viruses spread within and between human populations. The incorporation of BMDS within Bayesian phylogeography allows one to place dissimilarity data between species into a low-dimensional spatial representation while also considering their evolutionary dynamics from genetic data. Bedford et al., (2014) simultaneously characterize antigenic and genetic patterns of influenza by combining BMDS with an evolutionary diffusion process on the latent strain locations. They apply BMDS on hemmagglutination inhibition assay data to place the subtypes on a low-dimensional antigenic map. Holbrook et al., (2021) implement a similar Bayesian phylogenetic MDS model, but perform phylogeographic inference on pairwise distances arising from air traffic data. Additionally, Li et al., (2023) use phylogenetic BMDS on pairwise distances stemming from hepaciviruses to infer the viral locations in a lower dimensional geographic and host space.
Unfortunately, BMDS is difficult to scale to big data settings; computing the BMDS log-likelihood and gradient each have $\mathcal{O}(N^{2})$ complexity. Bedford et al., (2014) partially circumvent this problem by assuming that the observed data follow non-truncated Gaussian distributions, thereby avoiding the costly floating-point operations necessary to evaluate the Gaussian cumulative density function (CDF) in the truncated normal PDFs (2). However, there are benefits to using the truncated normal distribution: it appropriately accounts for non-negative dissimilarities, and its variance term is always less than that of its corresponding non-truncated normal distribution, resulting in more precise posterior inference. Holbrook et al., (2021) mitigate BMDS’s computational burden through massive parallelization using multi-core central processing units, vectorization and graphic processing units. They obtain substantial performance gains, but parallelization requires expensive hardware. In either case, these models still scale quadratically in the number of objects. We therefore develop a framework that reduces the time complexity to $\mathcal{O}(N)$ by inducing sparsity on the observed dissimilarity matrix. We perform experiments with simulated data and show that our sparse versions of BMDS (sBMDS) obtain significant speedups while preserving inferential accuracy. We then illustrate how one may use sBMDS within a larger hierarchical model by extending the types of phylogeographical models mentioned above under sparse assumptions, implementing sBMDS phylogenetic frameworks on air traffic data to analyze the geographic spread of four influenza subtypes. Additionally, we apply sBMDS to a collection of ArXiv paper abstracts and perform post-hoc clustering on posterior samples of the low-dimensional embeddings. We adopt a “bagged estimator” approach to propagate uncertainty quantification from sBMDS.
@@
In the following, we present two versions of sparse BMDS and prove that under simplistic conditions, the posterior latent locations are consistent for subsampled dissimilarity matrices (Section 2). In Section 3, we evaluate the empirical accuracy, sensitivity to model misspecification and computational performance of both methods. We apply sBMDS to the phylogeographic modeling of influenza variants and verify that we obtain similar migration rate estimates for both full and sparse BMDS models (Section 3.2). In Section 3.3, we apply sBMDS to a dataset representing ArXiv paper abstracts and recover the posterior probability that two manuscripts belong to the same subject-matter category. We conclude by summarizing our findings and discussing future research directions (Section 4).
@@ 2.1 Bayesian multidimensional scaling
Bayesian multidimensional scaling (BMDS) (Oh and Raftery,, 2001) models a set of $N$ objects’ locations as latent variables in low-dimensional space under the assumption that the observed dissimilarity measures follow a prescribed joint probability distribution. To set notation: for $A\subset\mathbb{R}$, let $N_{A}(\mu,\sigma^{2})$ denote the Gaussian distribution truncated to $A$; for $k\in\mathbb{N}$, let $[k]=\{1,2,\ldots,k\}$. Within BMDS, each observed dissimilarity measure $\delta_{nn^{\prime}}$ is the posited latent measure $\delta_{nn^{\prime}}^{*}$ plus a truncated Gaussian error:
$\delta_{nn^{\prime}}\sim N_{(0,\infty)}(\delta_{nn^{\prime}}^{*},\sigma^{2}),\ n\neq n^{\prime},\ n,n^{\prime}\in[N],$ (1)
@@
$\ell(\bm{\Delta},\sigma^{2})=-\frac{m}{2}\log(2\pi\sigma^{2})-\sum_{n<n^{\prime}}\biggr{[}\frac{(\delta_{nn^{\prime}}-\delta_{nn^{\prime}}^{*})^{2}}{2\sigma^{2}}+\log\Phi\biggr{(}\frac{\delta_{nn^{\prime}}^{*}}{\sigma}\biggr{)}\biggr{]},$ (2)
@@
$\nabla{\mathbf{x}_{n}}\ell(\bm{\Delta},\sigma^{2})=-\sum_{\begin{subarray}{c}n^{\prime}\in[N],\\ n^{\prime}\neq n\end{subarray}}\biggr{[}\biggr{(}\frac{(\delta_{nn^{\prime}}^{*}-\delta_{nn^{\prime}})}{\sigma^{2}}+\frac{\phi(\delta_{nn^{\prime}}^{*}/\sigma)}{\sigma\Phi(\delta_{nn^{\prime}}^{*}/\sigma)}\biggr{)}\ \frac{(\mathbf{x}_{n}-\mathbf{x}_{n^{\prime}})}{\delta_{nn^{\prime}}^{*}}\biggr{]}\equiv-\sum_{\begin{subarray}{c}n^{\prime}\in[N],\\ n^{\prime}\neq n\end{subarray}}\mathbf{r}_{nn^{\prime}}.$ (3)
@@ 2.2 Sparse likelihoods and their gradients
$\ell_{sp}(\bm{\Delta},\sigma^{2})=-\sum_{n=1}^{N}\sum_{\begin{subarray}{c}n^{\prime}\in J_{n,N},\\ n^{\prime}>n\end{subarray}}\biggr{[}\frac{1}{2}\log(2\pi\sigma^{2})+\frac{(\delta_{nn^{\prime}}-\delta_{nn^{\prime}}^{*})^{2}}{2\sigma^{2}}+\log\Phi\biggr{(}\frac{\delta_{nn^{\prime}}^{*}}{\sigma}\biggr{)}\biggr{]},$ (4)
@@
$\nabla{\mathbf{x}_{n}}\ell_{sp}(\bm{\Delta},\sigma^{2})=-\sum_{n^{\prime}\in J_{n,N}}\mathbf{r}_{nn^{\prime}}.$ (5)
We reduce the computational complexity of BMDS by including a small subset of couplings $J_{n,N}$ per object $n$, where $|J_{n,N}|\ll N$. Here, we discuss two possible strategies for choosing $J_{n,N}$. By a slight abuse of notation, we use $[a,b]$ to refer to a closed interval of either reals or integers, where the appropriate set should be obvious from context. The first option is to extract $B\in[N-1]$ off-diagonal bands of the observed dissimilarity matrix such that $J_{n,N}=[\max(1,n-B),\min(N,n+B)]\backslash\{n\}$ for all $n$. The second approach is to choose $L\in[N]$ objects called “landmarks” and select each landmark’s dissimilarities from the remaining $N-1$ objects, e.g., $J_{n,N}=[N]\backslash\{n\}$ for $n\in[L]$ and $J_{n,N}=[L]$ for $n\not\in[L]$. Essentially, this strategy retains a rectangular subset of the observed dissimilarity matrix by extracting $L$ columns (rows) of the data. We note there is no loss of generality in taking the first $n$ indices as landmarks rather than an arbitrary set because one can relabel the object indices without affecting the learned geometry. We refer to the first method as banded sBMDS (B-sBMDS) and the second method as landmark sBMDS (L-sBMDS). Alternative strategies for selecting index sets are possible, provided they satisfy Assumption 1 (Section 2.3). For instance, our B-sBMDS model assumes distances are measured on the real line and bands are defined as a contiguous interval. However, one could explore other forms of banded matrices, e.g., by selecting any set of entries in the distance matrix.
@@
Table 1: We extract the $(n,n^{\prime})$ pair from the off-diagonals of the observed and latent dissimilarity matrices for banded sBMDS (B-sBMDS) versus the columns for landmark sBMDS (L-sBMDS). $B/L$ refers to the number of bands ($B$) or landmarks ($L$), and the $+$ symbol indicates all couplings above are also included. The table on the right shows the calculated log-likelihoods as the number of bands/landmarks increases. Importantly, the bottom log-likelihoods are equal for both sBMDS variants and correspond to the full BMDS log-likelihood.
@@
Table 2: We extract the $(n,n^{\prime})$ pair from the observed and latent dissimilarity matrices to calculate the sBMDS gradients. On the left, the -- symbol, as in $(n,a-c)$, indicates pairs $(n,a);(n,b);(n,c)$. For example, pair $(3;4-5)$ means we include both pair $(3,4)$ and $(3,5)$. The $+$ symbol indicates all couplings to the left are also included, and [,][$\cdot$,$\cdot$][ ⋅ , ⋅ ] represents a vector. On the right is the gradient computed for each $\mathbf{x}_{n}$ of $\mathbf{X}$ as function of the number of bands/landmarks. Extracting the entire column of a landmark point gives the full BMDS gradient in $\mathbb{R}^{D}$ whereas banded sBMDS incrementally adds information to the row-wise gradients. Importantly, the rightmost gradients are equal for both sBMDS variants and correspond to the full BMDS gradient.
For B-sBMDS, the number of couplings is the number of elements in $B$ bands. The relationship between the number of bands and number of couplings $C$ is $C=\sum_{b=1}^{B}(N-b)$. We add one less coupling for each additional band. When the number of bands equals $N-1$, we return to the full BMDS case. Using a subset of the observed dissimilarity matrix reduces the burden of computing the BMDS likelihood and gradient to $\mathcal{O}(NB)$. Similar arguments hold for L-sBMDS, the likelihoods and gradients of which exhibit $\mathcal{O}(NL)$ time complexity.
For classical MDS, an analogous strategy to L-sBMDS already exists. In MDS, the rate limiting step is the calculation of the top $D$ eigenvalues and eigenvectors from a $N\times N$ matrix. Silva and Tenenbaum, (2004) propose applying classical MDS to $L$ landmark points, e.g., an $L\times N$ submatrix of the observed dissimilarity matrix, and then following a distance-based triangulation procedure to determine the remaining object coordinates. L-sBMDS uses the concept of randomly selecting $L$ landmarks as well, but integrates them into the BMDS framework, allowing inference on the entire model. Raftery et al., (2012) approximate the likelihood of their network data by taking a random subset of objects deemed to have no link, reducing the time complexity from $\mathcal{O}(N^{2})$ to $\mathcal{O}(N)$. In the context of a very different network model, they incorporate an array of covariates to model the probability of a link between objects $n$ and $n^{\prime}$ while our model is simpler, using no outside information to aid in determining locations in a latent space.